# Quantifying Emergence in Large Language Models

### v1

@@ 6 Limitations
We observe that LLMs exhibit a trend where larger model sizes correlate with higher emergence. This naturally prompts us to explore the empirical law between model parameters, dataset size, and emergence strength as Scaling Laws [3]. However, due to limited computational resources, we provide a preliminary study on smaller LMs and introduce a novel way of estimating the emergence of large LMs. Therefore, we would like to reiterate our contribution: pioneering the path for quantifying emergence in LLMs. Nonetheless, there still remains room for exploration such as a robust and efficient way of computing mutual information with higher dimensions and varied token lengths.
7 Conclusion
In this paper, we propose a quantification method to approximate the essential nature of the emergence of LLMs. Firstly, we provide a quantifiable definition of emergence and a resource-efficient algorithm to compute the strength. Secondly, under the settings of ICL and natural sentences, we conduct extensive experiments to validate the effectivenss of our algorithm, which not only corroborate existing findings, but also identify novel emergence patterns. Thirdly, implications for broader impacts are suggested such as the correlation between our quantification with hallucination and the feasibility to measure the emergence of closed-source LLMs.
@@ Appendix A Algorithm for Estimating Mutual Information
Appendix B Supplementary Materials for Finding 1
B.1 Limit Number of Shots
In Section 4, we expanded the 3 categories into input sequences containing 10 shots (20 tokens each). Table 3 illustrates the changes in emergence strength for each shot relative to its predecessor within these sequences. The emergence strength of 4 different LLMs generally approached their upper limits by the 6th and 7th shots. It is important to note that these results only indicate the existence of an upper limit to the contribution of shot quantity to emergence in ICL. They do not imply that the 6th and 7th shot universally represents the upper limit for all ICL tasks.
@@ B.1 Limit Number of Shots
B.2 Shot Length
To examine the emergence strength associated with shot lengths, we designed a shot format pertinent to sentiment analysis as follows:
@@ B.2 Shot Length
Figure 4 corroborates our hypothesis: within each shot, all tokens share a uniform emergence strength. This observation supports another intuitive viewpoint of ICL: an LLM gains greater confidence in the correctness of its predictions only when a new shot is introduced.
B.3 Cases of Inaccurate Generations with Excessive Shots
@@
Appendix C Supplementary Materials for Finding 2
measure t0 t1 t2 t3 t4 t5 t6 t7
baseline 4.69 4.69 9.46 9.37 15.32 15.02 28.44 29.47
model1 4.64 4.69 9.44 9.45 15.09 14.67 27.59 28.67
model2 4.64 4.68 9.44 9.28 15.27 14.66 44.08 36.37
model3 4.69 4.68 9.47 9.45 15.29 15.54 52.28 85.61
token1 2.82 2.83 6.88 6.85 11.08 10.95 16.83 16.09
token2 3.60 3.60 7.36 7.29 11.15 11.08 15.86 14.96
candidate 3.26 3.26 7.22 7.22 11.44 11.35 17.15 17.33
fusion1 3.84 3.84 7..63 7.64 11.66 11.49 17.45 16.28
fusion2 3.45 3.45 7.26 7.05 11.05 11.06 16.45 16.59
space 4.69 4.69 9.46 9.37 15.32 15.02 22.44 5.19
prefix 4.69 4.69 9.46 9.46 15.32 15.34 37.85 41.05
Table 4: “Ablation Study” of how emergence changes with different measures adopted. t0-t7 represent 1st - 8th token.
To investigate the influence of different factors on emergence strength, we treat the performance of GPT2-XL on the “country” dataset as a baseline and implemented a series of variations. First, we replace GPT2-XL with other LMs, namely model1 using GPT2-large, model2 using GEMMA, and model3 using OpenLlama. Second, we vary the dataset, forming data1 using “animal” dataset and data2 using “color” dataset. In addition, we use candidate to denote reduced candidates in the original “country” dataset (reducing the total number of countries from $25$ to $15$), and fusion1, fusion2 to denote mixed candidates where fusion 1 mixes data from “country” and “animal” domains, and fusion 2 mixes data from “country”, “animal”, and “color” domains. Last, we alter the input sequence, forming space by replacing the entity in the 4th shot with space+entity [^11] In the GPT-2 tokenizer, space+entity is treated as a new token.. prefix prepends a comma to the original first token.
As shown in Table 4, Differences in model emergence become apparent only when a sufficient number of shots are provided. Statistically, models with larger parameter sizes exhibit higher emergence. However, differences in data are apparent starting from the first shot, likely related to the domain of training data (token1, token2). Furthermore, depleting the diversity of shots effectively reduces the emergence effects in ICL (candidate). Lastly, the format of the prompt significantly influences emergence, explaining LLMs’ sensitivity to certain perturbations (space, prefix).
(a) GPT2-XL
(b) GEMMA
Appendix D Supplementary Materials for Finding 3
We observed variations in the emergence statistics of tokens at different positions within a sentence. Consequently, we systematically evaluated tokens at various positions within a sentence, as illustrated in Figure 5. Specifically, token0, token1, and token2 were derived from the same sample set A from OpenOrca, while token-3, token-2, and token-1 were taken from another sample set B from OpenOrca. Sample set A ensured that token0 was the initial token of the sentence, while set B ensured that token-1 was the last token of the sentence. This allowed us to measure differences in emergence statistics for tokens at the beginning, middle, and end of sentences across variable sentence lengths.
Figure 5 presents an interesting phenomenon: taking GPT2-XL and GEMMA as examples, GPT2-XL exhibits distinct responses to tokens at different sentence positions—emergence values increase at the beginning, stabilize in the middle, and rise again at the end. GEMMA, on the other hand, does not display such positional sensitivity. We hypothesize that this may be related to the different preprocessing methods used in the training data.
Appendix E Detailed Mutual Information Tables

### v2

@@ 8 Limitations
Position-wise Token: Given that mutual information intrinsically demands the distribution of two tokens to be valid, we require every token’s position to hold a specific meaning, such as representing the beginning or end of a sentence, the subject, predicate, and so forth. Hence, applying our estimator directly to existing tasks may result in a lack of interpretability as the token lengths and positions in the samples vary significantly.
Sample Amount: To ensure the accuracy of joint and marginal distributions of high-dimensional continuous representations, a tremendous number of samples is essential. We are attempting more mechanistic alternative methods, hoping to reduce sample size requirements in the future.
More Models and Tokens: It is evident that our experiments lack larger-sized models and analysis of long-length texts, especially for emergence and hallucination analysis. Given more computational resources, we would continue to expand these experiments.
@@ Appendix A Algorithm for Estimating Mutual Information
Appendix B Cases in Arithmetic Tasks
(a) GPT2-XL
(b) GEMMA
We have selected a total of 8 arithmetic tasks, as illustrated in the caption of Figure 3. For these tasks, we employed the 2-shots as the prompt templates for the ICL method. We randomly matched different shots for each sample. A representative example from each task is selected and presented as follows:
1 digit addition:
“What is 1 plus 0? A: 1, What is 4 plus 4? A: 8, What is 2 plus 7? A:”
1 digit division:
“What is 6 divided by 1? A: 6, What is 8 divided by 4? A: 2, What is 3 divided by 3? A:”
1 digit multiplication:
“What is 1 times 8? A: 8, What is 5 times 0? A: 0, What is 6 times 7? A:”
1 digit subtraction:
“What is 5 minus 2? A: 3, What is 7 minus 6? A: 1, What is 9 minus 0? A:”
2 digit addition:
“What is 53 plus 97? A: 150, What is 89 plus 25? A: 114, What is 75 plus 63? A:”
2 digit division:
“What is 72 divided by 9? A: 8, What is 81 divided by 27? A: 3, What is 18 divided by 3? A:”
2 digit multiplication:
“What is 95 times 55? A: 5225, What is 92 times 88? A: 8096, What is 43 times 42? A:”
2 digit subtraction:
“What is 25 minus 14? A: 11, What is 55 minus 36? A: 19, What is 80 minus 38? A:”
Appendix C Mutual Information in the Natural Scenario
We observed variations in the IE statistics of tokens at different positions within a sentence. Consequently, we systematically evaluated tokens at various positions within a sentence, as illustrated in Figure 5. Specifically, token0, token1, and token2 were derived from the same sample set A from OpenOrca, while token-3, token-2, and token-1 were taken from another sample set B from OpenOrca. Sample set A ensured that token0 was the initial token of the sentence, while set B ensured that token-1 was the last token of the sentence. This allowed us to measure differences in IE statistics for tokens at the beginning, middle, and end of sentences across variable sentence lengths.
Figure 5 presents an interesting phenomenon: taking GPT2-XL and GEMMA as examples, GPT2-XL exhibits distinct responses to tokens at different sentence positions—IE values increase at the beginning, stabilize in the middle, and rise again at the end. GEMMA, on the other hand, does not display such positional sensitivity. We hypothesize that this may be related to the different preprocessing methods used in the training data.
Appendix D Ablation Study for Semantics Sensitivity
measure t0 t1 t2 t3 t4 t5 t6 t7
baseline 4.69 4.69 9.46 9.37 15.32 15.02 28.44 29.47
model1 4.64 4.69 9.44 9.45 15.09 14.67 27.59 28.67
model2 4.64 4.68 9.44 9.28 15.27 14.66 44.08 36.37
model3 4.69 4.68 9.47 9.45 15.29 15.54 52.28 85.61
token1 2.82 2.83 6.88 6.85 11.08 10.95 16.83 16.09
token2 3.60 3.60 7.36 7.29 11.15 11.08 15.86 14.96
candidate 3.26 3.26 7.22 7.22 11.44 11.35 17.15 17.33
fusion1 3.84 3.84 7..63 7.64 11.66 11.49 17.45 16.28
fusion2 3.45 3.45 7.26 7.05 11.05 11.06 16.45 16.59
space 4.69 4.69 9.46 9.37 15.32 15.02 22.44 5.19
prefix 4.69 4.69 9.46 9.46 15.32 15.34 37.85 41.05
Table 3: “Ablation Study” of how IE value changes with different measures adopted. t0-t7 represent 1st - 8th token.
To investigate the influence of different factors on IE value, we treat the performance of GPT2-XL on the “country” dataset as a baseline and implemented a series of variations. First, we replace GPT2-XL with other LMs, namely model1 using GPT2-large, model2 using GEMMA, and model3 using OpenLlama. Second, we vary the dataset, forming data1 using “animal” dataset and data2 using “color” dataset. In addition, we use candidate to denote reduced candidates in the original “country” dataset (reducing the total number of countries from $25$ to $15$), and fusion1, fusion2 to denote mixed candidates where fusion 1 mixes data from “country” and “animal” domains, and fusion 2 mixes data from “country”, “animal”, and “color” domains. Last, we alter the input sequence, forming space by replacing the entity in the 4th shot with space+entity [^8] In the GPT-2 tokenizer, space+entity is treated as a new token.. prefix prepends a comma to the original first token.
As shown in Table 3, Differences in model IE become apparent only when a sufficient number of shots are provided. Statistically, models with larger parameter sizes exhibit higher IE. However, differences in data are apparent starting from the first shot, likely related to the domain of training data (token1, token2). Furthermore, depleting the diversity of shots effectively reduces the IE values in ICL (candidate). Lastly, the format of the prompt significantly influences IE, explaining LLMs’ sensitivity to certain perturbations (space, prefix).
Appendix E Supplementary Materials for Finding 1
E.1 Limit Number of Shots
In Section 4.1, we expanded the 3 categories into input sequences containing 10 shots (20 tokens each). Table 4 illustrates the changes in IE value for each shot relative to its predecessor within these sequences. The IE value of 4 different LLMs generally approached their upper limits by the 6th and 7th shots. It is important to note that these results only indicate the existence of an upper limit to the contribution of shot quantity to IE in ICL. They do not imply that the 6th and 7th shot universally represents the upper limit for all ICL tasks.
@@ E.1 Limit Number of Shots
E.2 Shot Length
To examine the IE value associated with shot lengths, we designed a shot format pertinent to sentiment analysis as follows:
@@ E.2 Shot Length
Figure 6 corroborates our hypothesis: within each shot, all tokens share a uniform IE value. This observation supports another intuitive viewpoint of ICL: an LLM gains greater confidence in the correctness of its predictions only when a new shot is introduced.
E.3 Cases of Inaccurate Generations with Excessive Shots
@@
Appendix F Detailed Mutual Information Tables